OpenStudy (anonymous):

Help please??? What is the equation of a circle with a center at (0, 0) and a radius of 4.5?

OpenStudy (anonymous):

Is it just that or is it written as (x-0)^2 + (y-0)^2=4.5^2?

OpenStudy (anonymous):

@10115658

OpenStudy (anonymous):

(x-0)^2 + (y-0)^2 is just x^2+y^2

OpenStudy (anonymous):

both are technically correct though x^2+y^2=4.5^2 looks nicer though

OpenStudy (anonymous):

and ignore my first answer lol